# Robust Multi-Agent Safety via Tube-Based Tightened Exponential Barrier Functions

### v1

@@ 4 Numerical Example
We demonstrate the proposed framework by simulating a leader-follower formation scenario in a two-dimensional space ($p=2$). The system consists of one leader and $N=5$ followers, each with distinct nonlinear dynamics and subject to external disturbances. Each agent’s dynamics are modeled with $n=3$ integrator blocks. The simulation is executed for a total of 30 seconds ($\text{sim}T=30$s), corresponding to $K_{\text{steps}}=300$ iterations with a sampling time of $T_{s}=0.1$s. The DMPC planner for each agent uses a prediction horizon of $H=5$. Two circular obstacles at $[1,1]$ and [1.5,0.5][-1.5,0.5] with radii $0.50$ and $0.65$ respectively include $0.15$ inflation margins. Ancillary feedback gains $K_{p}=-[15,4,15,8,6,8]$ were set via pole placement. Formation offsets are $\psi=[[-9;2],[-6;2],[0;2],[6;2],[9;2]]$ and eCBF gains $\kappa_{0}=30,\kappa_{1}=38,\kappa_{2}=3$. The leader starts at $[3,0,0,0,0,0]$ while followers begin at $1.0,-0.5,\text{zeros}(1,4)][$$0.8,0.4,\text{zeros}(1,4)]$,
@@
and $0.7,0.0,\text{zeros}(1,4)]$. Cost weights: $Q_{r}=50\,I_{10}$$P_{r}=10\,I_{10}$$R=0.01\,I_{2}$, $R_{du}=0.001\,I_{2}$. Formation coupling $\nu_{1}=\nu_{2}=1$ and stability parameters $\lambda_{1}=100,\lambda_{2}=50$. Figure 1 shows the information flow structure.
@@
Leader:
$\dot{x}_{1,1}^{0}=x_{2,1}^{0},\,\,\dot{x}_{1,2}^{0}=x_{2,2}^{0};\,\,\,\dot{x}_{2,1}^{0}=x_{3,1}^{0},\,\dot{x}_{2,2}^{0}=x_{3,2}^{0};$
$\dot{x}_{3,1}^{0}$ = $=-\,5\,\bigl(x_{3,1}^{0}\;+\;0.36\,x_{1,1}^{0}\;+\;0.84\,x_{2,1}^{0}\;$
+ $\quad+\;0.15\,(x_{1,1}^{0})^{3}\;-\;0.4\,\sin\bigl(0.8\,t)\bigr),$
$\dot{x}_{3,2}^{0}$ = $=-\,5\,\bigl(x_{3,2}^{0}\;+\;0.36\,x_{1,2}^{0}\;+\;0.84\,x_{2,2}^{0}\;$
+ $\quad+\;0.15\,(x_{1,2}^{0})^{3}\;-\;0.4\,\sin\bigl(0.8\,t+\frac{\pi}{2})\bigr).$
For each follower agent $i$, the common dynamics are:
$\dot{x}_{1,1}^{i}=x_{2,1}^{i},\,\,\dot{x}_{1,2}^{i}=x_{2,2}^{i};\,\,\,\,\dot{x}_{2,1}^{i}=x_{3,1}^{i},\,\,\dot{x}_{2,2}^{i}=x_{3,2}^{i};$
and the last block (jerk) is:
$\dot{x}_{3,1}^{i}$ $=f^{i}_{1}(x^{i})+u_{1}^{i}+w_{1}^{i}(t),$
$\dot{x}_{3,2}^{i}$ $=f^{i}_{2}(x^{i})+u_{2}^{i}+w_{2}^{i}(t),$
with $f^{i}_{1}$, $f^{i}_{2}$ and $w^{i}$ (disturbance) defined agent-wise below.
Follower $i=1$
$\hskip-11.0ptf^{1}_{1}(x^{1})=-k_{a}^{1}\!\Big(x_{3,1}^{1}+0.49\,x_{1,1}^{1}+1.12\,x_{2,1}^{1}+0.12\,(x_{1,1}^{1})^{3}\hfill\\ \hfill-0.25\,\tanh\!\big(0.6\,x_{1,1}^{1}\big)\Big),\\ f^{1}_{2}(x^{1})=-k_{a}^{1}\!\Big(x_{3,2}^{1}+0.49\,x_{1,2}^{1}+1.12\,x_{2,2}^{1}+0.12\,(x_{1,2}^{1})^{3}\hfill\\ \hfill-0.25\,\tanh\!\big(0.6\,x_{1,2}^{1}\big)\Big),\\ w_{1}^{1}(t)=0.20\,\sin(0.9\,t),\,w_{2}^{1}(t)=0.15\,\sin\!\Big(1.1\,t+\tfrac{\pi}{7}\Big).$
Follower $i=2$
$\hskip-11.0ptf^{2}_{1}(x^{2})=-k_{a}^{2}\!\Big(x_{3,1}^{2}+0.36\,x_{1,1}^{2}+0.84\,x_{2,1}^{2}+0.18\,(x_{1,1}^{2})^{3}\hfill\\ \hfill-0.15\,(x_{1,2}^{2})^{2}\,\tanh(x_{1,1}^{2})\Big),\\ f^{2}_{2}(x^{2})=-k_{a}^{2}\!\Big(x_{3,2}^{2}+0.36\,x_{1,2}^{2}+0.84\,x_{2,2}^{2}+0.18\,(x_{1,2}^{2})^{3}\hfill\\ \hfill-0.15\,(x_{1,1}^{2})^{2}\,\tanh(x_{1,2}^{2})\Big),\\ w_{1}^{2}(t)=0.18\,\sin(1.1\,t+0.3),\,w_{2}^{2}(t)=0.18\,\sin\!\Big(0.8\,t-\tfrac{\pi}{5}\Big).$
Follower $i=3$
$\hskip-11.0ptf^{3}_{1}(x^{3})=-k_{a}^{3}\!\Big(x_{3,1}^{3}+0.25\,\tanh(x_{1,1}^{3})+0.9\,\tanh(x_{2,1}^{3})\hfill\\ \hfill-0.20\,\sin(0.7\,t)\Big),\\ f^{3}_{2}(x^{3})=-k_{a}^{3}\!\Big(x_{3,2}^{3}+0.25\,\tanh(x_{1,2}^{3})+0.9\,\tanh(x_{2,2}^{3})\hfill\\ \hfill-0.20\,\sin\!\big(0.9\,t+\tfrac{\pi}{3}\big)\Big),\\ w_{1}^{3}(t)=0.18\,\sin(1.1\,t+0.3),\,w_{2}^{3}(t)=0.18\,\sin\!\Big(0.8\,t-\tfrac{\pi}{5}\Big).$
Follower $i=4$
$\hskip-11.0ptf^{4}_{1}(x^{4})=-k_{a}^{4}\!\Big(x_{3,1}^{4}+0.4225\,x_{1,1}^{4}+0.975\,x_{2,1}^{4}\hfill\\ \hfill+0.08\,\big(x_{1,1}^{4}+x_{1,2}^{4}\big)^{3}\Big),\\ f^{4}_{2}(x^{4})=-k_{a}^{4}\!\Big(x_{3,2}^{4}+0.4225\,x_{1,2}^{4}+0.975\,x_{2,2}^{4}\hfill\\ \hfill-0.08\,\big(x_{1,1}^{4}-x_{1,2}^{4}\big)^{3}\Big),\\ w_{1}^{4}(t)=0.12\,\sin(0.6t),\,w_{2}^{4}(t)=0.12\sin\!\big(0.6t+\tfrac{\pi}{2}\big).$
Follower $i=5$
$\hskip-11.0ptf^{5}_{1}(x^{5})=-k_{a}^{5}\!\Big(x_{3,1}^{5}+0.3025\,x_{1,1}^{5}+0.88\,x_{2,1}^{5}+0.15\,(x_{1,1}^{5})^{3}\hfill\\ \hfill-0.20\,\tanh\!\big(0.5\,x_{1,1}^{5}\big)\Big),\\ f^{5}_{2}(x^{5})=-k_{a}^{5}\!\Big(x_{3,2}^{5}+0.3025\,x_{1,2}^{5}+0.88\,x_{2,2}^{5}+0.15\,(x_{1,2}^{5})^{3}\hfill\\ \hfill-0.20\,\tanh\!\big(0.5\,x_{1,2}^{5}\big)\Big),\\ w_{1}^{5}(t)=0.22\,\sin\!\Big(0.9\,t+\tfrac{\pi}{8}\Big),\,w_{2}^{5}(t)=0.20\,\sin(1.0\,t).$
@@
Figure 2 shows the followers forming up with the leader while remaining outside the physical obstacle discs. Trajectories bend around the inflated safety envelopes (dashed cylinders); the closest passes occur near the right obstacle but stay outside its physical radius, so no violations occur. Pairwise separation is preserved - the eCBFs keep inter-agent distances above $d_{min}$. These results indicate that the collision- and obstacle-avoidance eCBFs are effective, and the tube tightening keeps the true trajectories within robust envelopes around the nominal plan.
@@ 5 Conclusion
This paper presented a distributed control framework that formally addresses the critical challenge of safe multi-agent navigation for high-order, nonlinear systems subject to disturbances. We introduced a unified architecture that synergistically combines the predictive capabilities of distributed Model Predictive Control (DMPC) with the formal safety guarantees of exponential Control Barrier Functions (eCBFs).
The core contribution is a method for robustly enforcing safety constraints on a nominal planner. By first constructing robust positively invariant (RPI) tubes to bound state errors, and then using their support functions to systematically tighten the eCBF constraints, the framework provides a formal margin that accounts for worst-case disturbances. This approach ensures that satisfying the tightened constraints on the nominal system guarantees forward invariance of safety sets for the true, disturbed system.
A key feature of this framework is its computational tractability. The synthesis method preserves the affine structure of the eCBF constraints relative to the control inputs, enabling the use of efficient, distributed quadratic programming. The resulting controller is therefore scalable and suitable for real-time implementation. As demonstrated in simulation, the framework achieves robust performance, maintaining bounded formation errors while successfully navigating a complex, obstacle-rich environment.
This research opens several promising avenues for future investigation. The framework could be extended to accommodate more dynamic and realistic communication scenarios, such as time-varying or switched network topologies and the presence of communication delays. Further research could explore the use of data-driven and learning-based methods to estimate disturbance bounds online, which would allow for the adaptive tuning of the RPI tubes to reduce conservatism. Finally, the integration of this provably-safe motion planning layer with higher-level decision-making algorithms would be a key step toward the deployment of fully autonomous and robust multi-agent teams in safety-critical applications.

### v2

@@ 4 Numerical Example
We validate the framework on a 2D leader–follower formation with one leader and $N=5$ followers, each modeled as a third-order $n=3$ nonlinear integrator-chain system with bounded disturbances. Simulations run for $30$ s with $T_{s}=0.1$ s $(300$ steps) and DMPC horizon $H=5$. Two circular obstacles are placed at $[1,1]$ and [1.5,0.5][-1.5,0.5] with radii $0.50$ and $0.65$ (plus $0.15$ inflation). Ancillary gains are $K_{p}=-[15,4,15,8,6,8]$ (pole placement), formation offsets are $\psi=[[-9;2],[-6;2],[0;2],[6;2],[9;2]]$, and eCBF gains are $\kappa_{0}=30,\kappa_{1}=38,\kappa_{2}=3$. The leader starts at $[3,0,0,0,0,0]$ and followers at $1.0,-0.5,\text{zeros}(1,4)][$$0.8,0.4,\text{zeros}(1,4)]$,
@@
and $0.7,0.0,\text{zeros}(1,4)]$. MPC weights are $Q_{r}=50I_{10},P_{r}=10I_{10},R=0.01I_{2},R_{du}=0.001I_{2}$, with coupling $\nu_{1}=\nu_{2}=1$ and stability parameters $\lambda_{1}=100,\lambda_{2}=50$. Figure 2 depicts the information-flow structure.
@@
Agent $\dot{x}_{1,k}^{i}=x_{2,k}^{i}$, $\dot{x}_{2,k}^{i}=x_{3,k}^{i}$, $k=1,2$, with $\dot{x}_{3,k}^{i}=f^{i}_{k}(x^{i})+u^{i}_{k}+w^{i}_{k}$ presented below:
[…]
Agent 1 $\begin{aligned} f^{1}_{1}(x^{1})=-k_{a}^{1}\!\big(x_{3,1}^{1}+0.49\,x_{1,1}^{1}+1.12\,x_{2,1}^{1}+0.12\,(x_{1,1}^{1})^{3}-0.25\,\tanh(0.6\,x_{1,1}^{1})\big),\\ f^{1}_{2}(x^{1})=-k_{a}^{1}\!\big(x_{3,2}^{1}+0.49\,x_{1,2}^{1}+1.12\,x_{2,2}^{1}+0.12\,(x_{1,2}^{1})^{3}-0.25\,\tanh(0.6\,x_{1,2}^{1})\big),\\ w_{1}^{1}(t)=0.20\,\sin(0.9\,t),\,w_{2}^{1}(t)=0.15\,\sin\!\big(1.1\,t+\tfrac{\pi}{7}\big).\par\end{aligned}$
Agent 2 $\begin{aligned} f^{2}_{1}(x^{2})=-k_{a}^{2}\!\big(x_{3,1}^{2}+0.36\,x_{1,1}^{2}+0.84\,x_{2,1}^{2}+0.18\,(x_{1,1}^{2})^{3}-0.15\,(x_{1,2}^{2})^{2}\,\tanh(x_{1,1}^{2})\big),\\ f^{2}_{2}(x^{2})=-k_{a}^{2}\!\big(x_{3,2}^{2}+0.36\,x_{1,2}^{2}+0.84\,x_{2,2}^{2}+0.18\,(x_{1,2}^{2})^{3}-0.15\,(x_{1,1}^{2})^{2}\,\tanh(x_{1,2}^{2})\big),\\ w_{1}^{2}(t)=0.18\,\sin(1.1\,t+0.3),\,w_{2}^{2}(t)=0.18\,\sin\!\big(0.8\,t-\tfrac{\pi}{5}\big).\end{aligned}$
Agent 3 $\begin{aligned} f^{3}_{1}(x^{3})=-k_{a}^{3}\!\big(x_{3,1}^{3}+0.25\,\tanh(x_{1,1}^{3})+0.9\,\tanh(x_{2,1}^{3})-0.20\,\sin(0.7\,t)\big),\\ f^{3}_{2}(x^{3})=-k_{a}^{3}\!\big(x_{3,2}^{3}+0.25\,\tanh(x_{1,2}^{3})+0.9\,\tanh(x_{2,2}^{3})-0.20\,\sin(0.9\,t+\tfrac{\pi}{3})\big),\\ w_{1}^{3}(t)=0.18\,\sin(1.1\,t+0.3),\,w_{2}^{3}(t)=0.18\,\sin\!\big(0.8\,t-\tfrac{\pi}{5}\big).\end{aligned}$
Agent 4 $\begin{aligned} f^{4}_{1}(x^{4})=-k_{a}^{4}\!\big(x_{3,1}^{4}+0.4225\,x_{1,1}^{4}+0.975\,x_{2,1}^{4}+0.08\,\big(x_{1,1}^{4}+x_{1,2}^{4}\big)^{3}\big),\\ f^{4}_{2}(x^{4})=-k_{a}^{4}\!\big(x_{3,2}^{4}+0.4225\,x_{1,2}^{4}+0.975\,x_{2,2}^{4}-0.08\,\big(x_{1,1}^{4}-x_{1,2}^{4}\big)^{3}\big),\\ w_{1}^{4}(t)=0.12\,\sin(0.6t),\,w_{2}^{4}(t)=0.12\sin\!\big(0.6t+\tfrac{\pi}{2}\big).\end{aligned}$
Agent 5 $\begin{aligned} f^{5}_{1}(x^{5})=-k_{a}^{5}\!\big(x_{3,1}^{5}+0.3025\,x_{1,1}^{5}+0.88\,x_{2,1}^{5}+0.15\,(x_{1,1}^{5})^{3}-0.20\,\tanh(0.5\,x_{1,1}^{5})\big),\\ f^{5}_{2}(x^{5})=-k_{a}^{5}\!\big(x_{3,2}^{5}+0.3025\,x_{1,2}^{5}+0.88\,x_{2,2}^{5}+0.15\,(x_{1,2}^{5})^{3}-0.20\,\tanh(0.5\,x_{1,2}^{5})\big),\\ w_{1}^{5}(t)=0.22\,\sin\!\big(0.9\,t+\tfrac{\pi}{8}\big),\,w_{2}^{5}(t)=0.20\,\sin(1.0\,t).\end{aligned}$
@@
Figure 3 shows the followers forming up with the leader while remaining outside the physical obstacle discs. Trajectories bend around the inflated safety envelopes (dashed cylinders); the closest passes occur near the right obstacle but stay outside its physical radius, so no violations occur. Pairwise separation is preserved – the eCBFs keep inter-agent distances above $d_{min}$. These results indicate that the collision- and obstacle-avoidance eCBFs are effective, and the tube tightening keeps the true trajectories within robust envelopes around the nominal plan.
@@ 5 Conclusion
This paper presented a robust safety framework for high-order nonlinear multi-agent systems under bounded disturbances, integrating tube-based error feedback with eCBF-constrained DMPC. The central technical contribution is a support-function-based tightening that converts per-agent RPI tubes into explicit scalar safety margins for pairwise and agent–obstacle eCBF constraints, so that nominal feasibility of the distributed planner implies forward invariance of the safety sets for the true disturbed trajectories. Crucially, the tightened constraints remain affine in the control input, preserving the tractability of the resulting distributed quadratic programs; simulations confirm safe formation navigation in obstacle-rich environments under persistent disturbances with bounded formation error. Future work includes extending the framework to time-varying and delayed communication topologies, reducing conservatism via online data-driven tube adaptation, and incorporating higher-level task planning for fully autonomous multi-agent deployment.